A group of chemists are conducting an experiment to produce a new liquid material. One chemical contains 15% sodium (Na) and the other chemical contains 30% sodium (Na). Once they mix the two samples the resulting chemical contains 22% sodium (Na). How many mili liters (ml) of each sample must be mixed to obtain 600 ml of the new chemical?
Answer by josmiceli(19441)

x = the chemical containing 15% Na
y = the chemical containing 30% Na
.15x + .30y = .22*600
also,
x + y = 600
y = 600 - x
.15x + .30(600 - x) = .22*600
.15x + 180 -.30x = 132
- .15x = -48
x = 320 ml
y = 600 - x
y = 280 ml
check
.15x + .30y = .22*600
.15*320 + .30*280 = .22*600
48 + 84 = 132
132 = 132
OK